OpenStudy (rebeccapink6398):

The area of a square is 225 square meters. What is the length of one side of this square? A. 15 m B. 112.5 m C. 75 m D. 50,625 m

OpenStudy (anonymous):

\(\sqrt{225}\)

OpenStudy (anonymous):

do you know how to find the area of a square?

OpenStudy (rebeccapink6398):

...no

OpenStudy (anonymous):

\[A = s*s\] \[A = s^2 \] \[225 = s^2 \] \[ sqrt (225) = s \] 15 = s

OpenStudy (rebeccapink6398):

thank you :D

OpenStudy (anonymous):

A.15